An asset acquired January 1, 2018, for $14,100 with an estimated 10-year life and no residual value is being depreciated in an equipment group asset account that has an average service life of eight years. The asset is sold on December 31, 2019, for $6,900. The entry to record the sale would be:

Multiple Choice

  • Cash 6,900
    Accumulated depreciation 3,525
    Loss on sale of equipment 3,675
    Equipment 14,100
  • Cash 6,900
    Loss on sale of equipment 7,200
    Equipment 14,100
  • Cash 6,900
    Equipment 6,900
  • Cash 6,900
    Accumulated depreciation 7,200
    Equipment 14,100
Answer #1

Cost of assets = $14,100

Date of purchase = January 1, 2018

Useful life = 8 years

Date of sale = December 31, 2019

Sale price = $6,900

Equipment has been used for 2 years

Annual depreciation = Cost of equipment / Useful life

= 14,100/8

= $1,762.50

Accumulated depreciation for 2 years = Annual depreciation x 2

= 1,762.5 x 2

= $3,525

Book value equipment on the date of sale = Cost of assets - Accumulated depreciation for 2 years

= 14,100-3,525

= $10,575

Loss on sale of equipment = Book value equipment on the date of sale - Sale price

= 10,575-6,900

= $3,675

The entry to record sale would be:

  • Cash 6,900
    Accumulated depreciation 3,525
    Loss on sale of equipment 3,675
    Equipment 14,100

1st option is correct option.

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.

Thanks!!!